Fermi National Laboratory


Accelerator Update

Friday March 19
The day shift began with the fifth day of shutdown. There was beam in Linac and Booster with Booster experts tuning.

At 8:14 AM, the Antiproton Source (Pbar) feeder outage ended.

Linac experts halted beam to make an access at 11:46 AM. They were out of access an hour later.

Operations reestablished beam to Booster at 1:13 PM.

At 2:24 PM, the vacuum at Tevatron (TeV) sector E4 went bad while technicians were working on the system. A vacuum valve failed.

Operations performed a Main Injector (MI) highpot at 11:57 PM. The test showed a ground fault somewhere between MI-10 and MI-50. The search for this fault will occur later in the morning.

RF techs brought a Booster RF station (BRF16) back on line, from an earlier problem, around 5:30 PM, but there was no beam to Booster because of a Linac intertank phase problem between Drift Tubes LRF1 and LRF2.

BRF16 tripped off four more times over a one and half hour period starting at 6:36 PM.

A Linac tech began work on the LRF2 phase problem at 7:55 PM.

Cryo system techs at CHL began the cryo system cooldown of TeV sector E4 at 10:41 PM.

Saturday March 20
The midnight shift began the sixth and last day, of shut down for all machines except the TeV. TeV work in sectors A4, E4, and B4 continued. Booster experts conducted studies.

At 12:05 AM, BRF16 tripped off and then tripped again twenty minutes later. It reset both times.

Booster studies ended at 1:11 AM.

Operators completed their interlock of Pbar at 3 AM.

Operators and a Recycler expert accessed MI to search for the ground fault at 4:17 AM.

At 5:14 AM, vacuum techs at TeV sector A4 reported that they would be finished with their work and out of the tunnel by 6 AM.

BRF16 tripped off at 7:21 AM. It reset okay.

Cryo system techs began work on the TeV sector F3 system at 8:17 AM.

Operations noted that the LRF2 phase problem returned at 9:51 AM. Operators called in a Linac tech and he had Linac back in operation by 11 AM.

Cryo system techs began the TeV sector E2 cooldown at 11:37 AM.

At 1:13 PM, cryo system techs began warming the TeV sector B3 system for alignment work.

Pbar startup began at 2:30 PM.

By 5:47 PM, operators had interlocked TeV sector F and the MI tunnels.

Operations established beam to the MI dump at 8 PM, but was seeing beam closure problems.

At 10 PM, CHL techs were cooling down the TeV sectors A4, B4, and C2 cryo systems.

Operators called in an instrumentation tech at 10:10 PM to investigate a MI BPM problem. MI studies, which were to begin at 11:48 PM, were delayed due to the BPM problem.

Sunday March 21
The midnight shift began with MI BPM repairs and TeV work continuing.

CHL began warming the cryo system at TeV sector A3 at 12:23 AM.

Operations reported that there were still BPM problems at MI-8 at 1:23 AM.

MI studies began at 3:30 AM and finished around 7:30 AM.

At 7:43 AM, LRF2 tripped off for the first of two times in about ten minutes.

Operations reported at 8 AM that a Debuncher kicker (D:IKIK) had tripped off and wouldn't reset. Operators contacted an expert and then got the kicker back up by 8:24 AM.

Operations reported having a problem getting MI beam to the P1 transfer line at 8:34 AM.

LRF2 tripped off at 9:11 AM and then again twenty minutes later.

At 11:38 AM, CHL began the TeV sector A3 cryo system cooldown.

At noon, experts looked at the MI pictures operators had taken of a major bend magnet (V701) and determined it was wired wrong. The experts access MI at 12:14 PM to make repairs. They were out of access by 1:06 PM. They reported finding a buss short and reversed cables. Experts completed another access at 3:24 PM to reverse polarities on other V701 elements. Here are pictures that show what they found and corrected.




At 4 PM, experts and operators got beam through MI and to the Pbar dump at AP0.

Monday March 22
The midnight shift began with Pbar experts conducting studies, and with work in the TeV continuing.

At 1:24 AM, Pbar studies reported having a problem getting beam into the Accumulator. They called in a tech to look at a septum (A:ISEPV2).

Operations prepared Pbar for access at 1:32 AM. The experts accessed, made repairs, came out, and had Pbar stacking by 2:45 AM.

A new round of Pbar studies began at 4:38 AM.

Operations established a proton stash in the Recycler at 4:49 AM.

Operators had MiniBooNE interlocked by 6:55 AM.

Today's Plans
The plan for today is to continue TeV work while the other machines conduct studies. Beam to MiniBooNE later this morning for testing devices.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 3/22/2004   email Fermilab